# veScale-FSDP: Flexible and High-Performance FSDP at Scale

### v1

@@ 4 RaggedShard for Flexibility
Existing sharding formats.   The second format is the Row-wise (Even) Shard, where a tensor is evenly partitioned along a dimension, with equal-sized shards assigned to each device. This design improves flexibility by enabling non-element-wise computations on sharded tensors and allowing dimension redistribution via All2All collectives. However, it still faces challenges with block-wise quantization, as evenly divided shards are not guaranteed to align with block boundaries. This row-wise sharding format serves as the foundation of FSDP2.
@@
To support combinations of multiple parallelization strategies, RaggedShard needs to compose cleanly with existing DTensor placements. RaggedShard is orthogonal to both replicated and partial-value placements and veScale-FSDP specially handles the Shard placement. In practice, TP uses Shard(0) and Shard(1) for column- and row-wise tensor parallelism; EP can be encoded as Shard(0) along the expert dimension. By convention, EP/TP is applied before FSDP. In PyTorch, however, the DTensor placement list is organized in the opposite order of conceptual application (see Figure 5): a tensor shown with placements (RaggedShard, Shard(0)) is partitioned as Shard(0) followed by RaggedShard. veScale-FSDP reconciles this by: (i) for Shard(0), introducing a dedicated placement StridedRaggedShard that carries reordering/stride metadata and performs Meanwhile, RaggedShard, as an extended DTensor placement, can offers the checkpointing capability by directly reusing DTensor-based checkpointing stacks (e.g., PyTorch Distributed Checkpoint [dcp]) for failure recovery and also inheriting their optimizations such as communication-free sharded checkpointing.

### v2

# arXiv:2602.22437v2[cs.DC] 27 Feb 2026

@@ 4 RaggedShard for Flexibility
Existing sharding formats.   Figure 4 summarizes existing sharding formats for FSDP tensors. The most common yet conventional format is the Element-wise Shard, where tensors are partitioned arbitrarily across devices without respecting structural boundaries. This leads to dangling elements on each device and a loss of tensor shape and stride information. Consequently, this format severely limits flexibility: (i) non-element-wise operations such as matrix multiplications break on misaligned elements, (ii) collective communication for redistributing tensors across dimensions becomes complex and inefficient, and (iii) quantization methods such as block-wise FP8 quantization cannot align block sizes with random shard boundaries, introducing substantial communication overhead for cross-boundary synchronization. Unfortunately, this element-wise sharding design remains the backbone of DeepSpeed and FSDP1.
The second format is the Row-wise (Even) Shard, where a tensor is evenly partitioned along a dimension, with equal-sized shards assigned to each device. This design improves flexibility by enabling non-element-wise computations on sharded tensors and allowing dimension redistribution via All2All collectives. However, it still faces challenges with block-wise quantization, as evenly divided shards are not guaranteed to align with block boundaries. This row-wise sharding format serves as the foundation of FSDP2.
@@
To support combinations of multiple parallelization strategies, RaggedShard needs to compose cleanly with existing DTensor placements. RaggedShard is orthogonal to both replicated and partial-value placements and veScale-FSDP specially handles the Shard placement. In practice, TP uses Shard(0) and Shard(1) for column- and row-wise tensor parallelism; EP can be encoded as Shard(0) along the expert dimension. By convention, EP/TP is applied before FSDP. In PyTorch, however, the DTensor placement list is organized in the opposite order of conceptual application (see Figure 5): a tensor shown with placements (RaggedShard, Shard(0)) is partitioned as Shard(0) followed by RaggedShard. veScale-FSDP reconciles this by: (i) for Shard(0), introducing a dedicated placement StridedRaggedShard that carries reordering/stride metadata and performs the required reshuffle when materializing the full tensor; and (ii) for Shard(dim>0), adapting the ragged sharding granularity so it never cuts into that dimension, choosing the granularity as Least Common Multiple (LCM): LCM(stride(dim), user-defined granularity).
Meanwhile, RaggedShard, as an extended DTensor placement, can offers the checkpointing capability by directly reusing DTensor-based checkpointing stacks (e.g., PyTorch Distributed Checkpoint [dcp]) for failure recovery and also inheriting their optimizations such as communication-free sharded checkpointing.